Starlike and Close-to-Convex Functions with Respect to Symmetric-Conjugate Points

Hassoon Al–Amiri, Bowling Green, Dan Coman, Ann Arbor, U.S.A. and Petru T. Mocanu, Cluj–Napoco, Romania

Keywords: Holomorphic functions, starlike and close-to-convex functions, symmetric-conjugate points, differential subordinations AMS Subject classification: 30C45

Abstract: In this paper four new classes of close-to-convex functions are introduced by combining the concept of starlike functions with respect to symmetric points with that of starlike (close-to-convex) functions with respect to conjugate points. Using differential subordinations, certain results concerning inclusion relations and integral operators defined on these subclasses are obtained.